find the slope of the line that passes through each pair of points.
1) (7/3, 4/3)(-1/3, 2/3)

Answer by nycsharkman(136)

The slope is rise/run.
I will use letters to represent your points.
(a, b) (c, d) = the two given points.
slope = (d minus b) divided by (c - a)
slope = (2/3 - 4/3) divided by (-1/3 - 7/3)
slope = -2/3 divided by -8/3
slope = 1/4
Got it?